# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

### v3

@@ 4 Methodology
In Section 2, we reviewed prior studies related to traffic control on inducing robustness and resilience using different RL algorithms. Here, we demonstrate how we integrate the concept of antifragility and enhance the performance of traffic control algorithms.
@@ 4.1 RL algorithm
In RL algorithms, an agent or multiple agents interact with a preset environment and improve the performance of decision-making, defined as action $a_{t}$ in an action space $\mathcal{A}$, based on the observable state $s_{t}$ in the state space $\mathcal{S}$ and the reward, defined as $r_{t}=\mathcal{R}(s_{t},a_{t})$, where $\mathcal{R}$ is the reward function. The improvement of decision-making is commonly realized through a deep neural network as the function approximator. The RL algorithm applied in this work is Deep Deterministic Policy Gradient (DDPG), as proposed in Lillicrap et al. (2015). By applying an actor-critic scheme, DDPG can manage a continuous action space instead of a discrete action space. Furthermore, Zhou and Gayah (2021) has demonstrated that an RL algorithm with a continuous action space can achieve superior performance. The DDPG algorithm can be divided into two main components, namely the actor and the critic, which are updated at each step through policy gradient and Q-value, respectively. The scheme of the DDPG algorithm applied in this paper is schematically illustrated in Fig. 2.
@@
Afterward, the actor network can be updated with sampled deterministic policy gradient:
@@ 4.2 Antifragility and the antifragile terms in RL
As reviewed in Section. 2.2, robustness and resilience can be induced by integrating proper information into the RL algorithms. Aligning with the same idea, we incorporate antifragile derivatives and redundancy terms based on the methodologies reviewed in Section 2.3. First, the state $s_{t}\in\mathcal{S}$ is defined distinctively in different methods applied in this work. The benchmark work in Zhou and Gayah (2021) establishes the state as $s_{t}=[n_{ij}(t),q_{ij}(t)]$, with $q_{ij}(t)$ regarded as an estimate of average daily traffic demand. However, real-time traffic demand $q_{ij}(t)$ can hardly be acquirable, particularly under an environment with disruptions. Therefore, we propose a state representation incorporating the derivatives of vehicle accumulation, which can be computed in the discrete form as:
$s_{t}=[n_{ij}(t),\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ (8)
@@
The action $a_{t}\in\mathcal{A}$ is defined the same as the control variables $u_{ij}(t)$. For the reward $r_{t}$, while Zhou and Gayah (2021) uses merely the completion rate, in our proposed algorithm, the reward is defined with additional two terms, i.e., the damping term $r_{dam}(t)$ and the redundancy term $r_{red}(t)$, as Eq. 4 shows. When $\Delta n_{ij}(t)$ and $\Delta^{2}n_{ij}(t)$ are incorporated into the state $\mathcal{S}$ of the algorithm, significant oscillations in the perimeter control variables can often be observed, especially under scenarios with disruptions. As perimeter control is composed of coordinated traffic lights at the border of different regions, oscillating actions will result in fast varying green splits between consecutive cycles. Even though such oscillations may have little impact on certain performance indicators such as trip completion, the operation of traffic lights in the real world should be as stable as possible. Therefore, a damping term $r_{dam}(t)$ is introduced into the reward function to penalize potential oscillatory actions. The exponent $\xi$ determines how fast the penalty decays when the change of control variables becomes small, indicating a large $\xi$ only penalizes the agent when the oscillation is substantial.
$r_{dam}(t)=-\sum_{i,j\in\{1,2\}}|u_{i,j}(t)-u_{i,j}(t-1)|^{\xi},\quad i\neq j$ (9)
@@
The last term $r_{red}(t)$ in the objective function $J_{anti}$ acts as an additional term to build up redundancy in the system, resembling the buffer times in train scheduling. However, instead of a fixed buffer time commonly seen among most European railway networks (Jovanović et al., 2017), we create a dynamic redundancy term emphasizing the derivates similar to $\Delta n_{ij}(t)$ and $\Delta^{2}n_{ij}(t)$ in the state $\mathcal{S}$. Here, we summarize $r_{red}(t)$ as the sum of two terms $r_{red}(t)=H(t)+\Delta H(t)$, with $H(t)$ being an overall term representing the first derivative and $\Delta H(t)$ representing the second derivative, which can be further expanded as:
@@
The binary variable $\alpha_{i}(t)$ in $H(t)$ was designed to reward the agent when moving towards the desired direction on the MFD. For instance, when the traffic state lies in the congested zone of the MFD, the gradient of any step will be negative. However, a penalty should be applied when vehicle accumulation gets larger, while rewarding the agent accumulation decreases.
@@
$f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})$ computes a reduction factor to constrain the impact of the $r_{red}(t)$ term when the accumulation around the critical accumulation $n_{i,\rm crit}$, where the $r_{red}(t)$ term should have the greatest impact. We use a modified trigonometric function to realize this purpose. Other functions, such as normal distribution, are also valid for achieving the same purpose.
@@
We illustrate the impact of $H(t)$ and $\Delta H(t)$ on building redundancy on the MFD in Fig. 3 and 4. The first derivative $H(t)$ in Fig. 3(a) rewards the agent when moving toward the critical accumulation $n_{i,\rm crit}$ to maximize trip completion. However, when $n$ approaches $n_{i,\rm crit}$, this term drops significantly and becomes a penalty when $n$ exceeds $n_{i,\rm crit}$. In Fig. 3, we showcase the influence of this term on the MFD. With increasing weight coefficient $\omega_{h}$, $n_{i,\rm crit}$ of the modified MFD becomes lower compared to the original, and the reward decreases faster after the accumulation exceeds $n_{i,\rm crit}$. In this way, redundant overcompensation has been established to prevent accumulation from exceeding the critical accumulation when disruption happens unexpectedly.
@@
4.3 Model Predictive Control
Model Predictive Control (MPC) is a widely recognized control approach extensively used for regulating dynamic systems across various engineering fields, particularly in transportation and perimeter control. Readers interested in Model Predictive Control (MPC) and its applications in traffic engineering can refer to Geroliminis et al. (2013); Haddad and Mirkin (2017). The MPC toolkit applied in this paper is introduced in Lucia et al. (2017), which uses the CasADi framework (Andersson et al., 2019) and the NLP solver IPOPT (Wächter and Biegler, 2006).
@@ 5 Experiment application
This section introduces the most important settings for the numerical simulation to showcase antifragile perimeter control and the corresponding performance evaluation. In addition to No Control (NC), we investigated two state-of-the-art perimeter control algorithms compared to our proposed antifragile algorithm, which are summarized below:
    NC where the system operates without intervention.
    MPC proposed in Geroliminis et al. (2013).
    State-of-the-art RL-based perimeter control algorithm proposed in Zhou and Gayah (2021) as the RL benchmark.
The performance of the traffic control algorithms is studied under two different types of disruptions as illustrated in Fig. 5, i.e., a demand disruption like surging traffic demand, or a supply disruption, such as the change of MFD for a specific network due to adversarial weather or major accidents. Furthermore, to showcase both proto-antifragility and progressive antifragility, this work examines the impact of different manifestations of disruption magnitudes. In the short term, the magnitudes are assumed to remain relatively stable with minor stochastic variations. In the long term, however, as traffic demand continues to increase and contributes to more frequent traffic incidents, the magnitude is also expected to grow, which can be subject to real-world uncertainties as well. The static and incremental magnitude of disruptions is intended for testing proto- and progressive antifragility. To sum up, the combination of 1) demand and supply disruptions, 2) static and incremental disruptions, and 3) with and without uncertainty in disruptions has resulted in a total number of 8 scenarios.
@@ 5.1 Simulation environment parametrization
We simulate a cordon-shaped urban network with inner and outer regions represented by different MFDs, as Fig. 1(b) shows. These MFDs are largely the same as in Zhou and Gayah (2021) and were originally approximated through the Yokohama loop detector dataset (Geroliminis and Daganzo, 2008). However, since the MFDs in Zhou and Gayah (2021) are formulated as piecewise functions and are, although continuous, not differentiable, and indifferentiability within MFDs can cause fluctuations when computing the first and second differences and harm the efficacy of the redundancy term. Therefore, a minor modification has been made to slightly increase the gridlock accumulation by $3\%$, so that the whole MFD can be both continuous and differentiable. Other important indicators, e.g., critical accumulation and maximal trip completion, remain the same as in Zhou and Gayah (2021).
Even though real-world disruptions can happen to both the outer and the inner regions and exert a negative influence, we focus on the bottleneck scenarios where the potential of perimeter control can be better reflected. Therefore, for a demand disruption, a surging traffic demand is assumed to be generated from the outer region toward the inner region, whereas for a supply disruption, the inner MFD is lowered as a result of disruptive events while the outer MFD remains intact. Demand and supply disruptions also demonstrate different training difficulties for the two RL-based algorithms involved, which will be detailed in Section 6.
Traffic demand under no disruption is approximated based on Geroliminis and Daganzo (2008). As the real-world peak hour traffic demand profile has more resemblance to a Gaussian distribution instead of being a simple trapezoidal shape (Mazloumi et al., 2010), the base demand $q_{ij}(t)=[q_{11}(t),q_{12}(t),q_{21}(t),q_{22}(t)]$ is illustrated as the blue curves in Fig. 6(a). It consists of two components, which are a constant value $q_{ij,c}=[q_{11,c},q_{12,c},q_{21,c},q_{22,c}]=[0.3,0.35,0.2,0.25]$ in $\mathrm{veh/s}$ and a Gaussian term $q_{ij,N}(t)=C_{ij,N}\mathcal{N}_{ij}(\mu_{ij,N},\sigma_{ij,N})$ with total number of vehicles following Gaussian as $C_{ij,N}=[3000,15000,2000,4000]$ and the mean and variance of the distribution being $\mu_{ij,N}=[1800,1800,1800,1800]$ and $\sigma_{ij,N}=[1200,1500,900,1200]$ in seconds. Under a demand disruption scenario, $q_{12,N}$ is increased to simulate traffic flowing from the periphery into the city center. For testing progressive antifragility, a compounding factor of $0.018$ per episode is applied to escalate $q_{12,N}$, with the maximal demand disruption after 25 episodes reaching $(1.018^{25}-1)C_{12,N}\mathcal{N}_{12}(\mu_{12,N},\sigma_{12,N})$. On the other hand, the magnitude for validating proto-antifragility is set to be static as $(1.018^{25/2}-1)C_{12,N}\mathcal{N}_{12}(\mu_{12,N},\sigma_{12,N})$.
(a) Base demand profile and disruptions
(b) MFD and disruptions for the inner region
@@
Supply disruptions can manifest in various forms, such as adverse weather conditions, traffic accidents, and road maintenance. However, limited research has been dedicated to understanding the exact correlation between such events and their impact on the MFD. Recent work in Ambühl et al. (2020) proposed the concept of infrastructure potential $\lambda$ to represent how efficiently the network infrastructure is used considering the performance loss due to vehicle infrastructure interaction, and a smaller $\lambda$ indicates the infrastructure is utilized with higher efficiency. In this research, a $\lambda$ equal to $0.07$ is chosen for the MFD without disruption representing the inner region, which lies in the common range of $\lambda$ based on data from real-world cities (Ambühl et al., 2020). And for the incremental magnitude of supply disruptions to test progressive antifragility, $\lambda$ is escalated by $0.003$ per episode with the maximal $\lambda$ after 25 episodes being $0.145$, as the solid red curve shows in Fig. 6(b). For static supply disruptions to test proto-antifragility, the magnitude of disruption is set to be half of the maximal disruption. Note that since the polynomial MFD in Geroliminis and Daganzo (2008) and the MFD approximated with $\lambda$ are essentially two different functional forms, there are still marginal deviations between each other after careful parameter calibration, as illustrated with the blue and gray curves in Fig. 6(b).
The total simulation duration $T$ is 3 hours with each time step $\Delta t$ as $180$ seconds, and the third hour has little traffic demand and mainly serves as the unloading process. Another notable constraint is the lower and upper bounds for the perimeter control variable $u_{ij}(t)\in[0.1,0.9]$ to represent the real-world constraints from traffic signals (Geroliminis et al., 2013). The initial vehicle accumulation is set to be $n_{ij}(0)=[600,1300,300,2400]$ so that the accumulation remains approximately at an equilibrium at the beginning of the simulation.
@@
Each scenario is run $25$ times since randomness is inherent in RL, and its performance may vary across simulation iterations. Each iteration lasts for $75$ episodes, with the first $50$ under no additional disruption so that the RL agent can be properly trained under the base demand profile with predefined training noise. Either static or incremental magnitude of disruptions is introduced in the subsequent $25$ episodes, with the value deliberately set to be the same as the number of simulation iterations. Under disruption uncertainties, as illustrated by the zigzag lines for the simulation scenarios in Fig. 5. A list of multipliers $\epsilon=[\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{24},\epsilon_{25}]$ following a normal distribution $\mathcal{N}\sim(1,0.15)$ is randomly generated to introduce uncertainties by computing the dot product with the list of disruption magnitudes. The multiplier list will be shuffled by 1 for each simulation iteration. For example, the list would be $\epsilon=[\epsilon_{2},\epsilon_{3},\ldots,\epsilon_{25},\epsilon_{1}]$ for the second iteration. Through reshuffling, the disruption magnitudes from different episodes will experience exactly the same list of uncertainty multipliers but with different sequences.
@@
The most important hyperparameters for both the baseline and the proposed RL algorithms should be the same and are summarized in Table. 2 below. Note that the minimal learning rates and noise scale are not set to be an extremely small value as common RL algorithms do, we aim to get a trade-off between optimality and adaptiveness when the algorithm is experiencing disruptions, which can be demonstrated by the results for both good performance and antifragility in Section 6. The exponent coefficient $\xi$ for the damping term $r_{dam}(t)$ in Eq. 9 to counter the potential oscillations actions is set to be $6$. The weights $\omega_{h}$ and $\omega_{\Delta h}$ for the redundancy term $r_{red}(t)$ is set to be $0.1$ and $0.2$. The number of simulations generated under the Ape-X architecture is 32 per training episode.
@@
Learning rate decay 0.975 0.975

### v4

@@ 4 Methodology
In this section, we first discuss the methodologies of the benchmarks and use their benefits and drawbacks to motivate the development of our proposed antifragile perimeter control algorithm.
4.1 Model predictive control and moving horizon estimation
Model Predictive Control (MPC) is a widely recognized control approach extensively used for regulating dynamic systems across various engineering fields, particularly in transportation and perimeter control. Readers interested in MPC and its applications in traffic engineering can refer to Geroliminis et al. (2013); Haddad and Mirkin (2017). The MPC toolkit applied in this paper is introduced in Lucia et al. (2017), which uses the CasADi framework (Andersson et al., 2019) and the NLP solver IPOPT (Wächter and Biegler, 2006).
In Sirmatel and Geroliminis (2020), Moving Horizon Estimation (MHE) is introduced to form an MPC-MHE framework for perimeter control. MHE enables real-time estimation of internal system states by solving an optimization problem that minimizes the discrepancy between predicted outputs and noisy sensor measurements. It further addresses the data observability issues in Geroliminis et al. (2013) as illustrated in Fig. 1, that is, the information on both the OD accumulation $n_{ij}(t)$ and the newly generated demand $q_{ij}(t)$ is almost impossible to acquire in real-time. Therefore, four different combinations of data availability have been tested, i.e., one from either $n_{ij}(t)$ or $[n_{i}(t),M_{ij}(t)]$ and another one from either $q_{ij}(t)$ or $q_{i}(t)$. The accumulation $n_{i}(t)$ for region $i$ can be approximated with loop detector measurements, while interregional trip completion $M_{ij}(t)$ can be directly counted through detectors at the intersections forming the perimeter border. The authors also claim $q_{i}(t)$ can be available with some applications, such as a substantial percentage of GPS information being readily accessible. However, due to the nature of the prediction horizon in MPC-MHE, such data has to be made available in both real-time and the near future, especially under disruptive conditions, rendering it impossible to collect. We will further tackle such observability issues with the proposed antifragile perimeter control algorithm in the following Section 4.3.
@@ 4.2 RL algorithm
In RL algorithms, an agent or multiple agents interact with a preset environment and improve its decision-making capacilities, defined as action $a_{t}$ in an action space $\mathcal{A}$, based on the observable state $s_{t}$ in the state space $\mathcal{S}$ and the reward, defined as $r_{t}=\mathcal{R}(s_{t},a_{t})$, where $\mathcal{R}$ is the reward function. The improvement of decision-making is commonly realized through a deep neural network as a function approximator. The RL algorithm applied in this work is Deep Deterministic Policy Gradient (DDPG), as proposed in Lillicrap et al. (2015). By applying an actor-critic scheme, DDPG can manage a continuous action space instead of a discrete action space. Furthermore, Zhou and Gayah (2021) has demonstrated that an RL algorithm with a continuous action space can achieve superior performance. The DDPG algorithm can be divided into two main components, namely the actor and the critic, which are updated at each step through policy gradient and Q-value, respectively. The scheme of the DDPG algorithm applied in this paper is schematically illustrated in Fig. 3.
@@
Afterward, the actor network can be updated with the sampled deterministic policy gradient:
@@ 4.3 Antifragility and the antifragile terms in RL
As reviewed in Section. 2.2, proper information can be integrated into the RL algorithms to induce robustness and resilience. Following the same idea, we incorporate antifragile derivatives and redundancy terms based on the methodologies reviewed in Section 2.3. First, the benchmark work in Zhou and Gayah (2021) establishes the state as $s_{t}=[n_{ij}(t),q_{ij}(t)]$, with $q_{ij}(t)$ regarded as an estimate of average daily traffic demand. However, real-time traffic demand $q_{ij}(t)$ can hardly be acquired, particularly in an environment with disruptions. Therefore, we propose a state representation incorporating both the first and second order derivatives of vehicle accumulation, which can be computed as the first and second order differences between the vehicle accumulation $[\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ between two consecutive time steps in the discrete format. Similar to Sirmatel and Geroliminis (2020) discussed in Section 4.1, we also tackle the data observability issue in this work, by integrating trip completion $M_{ij}(t)$ and replacing OD-pair accumulation $n_{ij}(t)$ with simply regional accumulation $n_{i}(t)$ so as its first and second differences. It should be noted that $M_{ij}(t)$ in Sirmatel and Geroliminis (2020) refers specifically to the interregional transfer flow $M_{ij}(t),(i\neq j)$. However, when considering a rather homogenous average trip length, the regional trip completion rate $M_{i}(t)$ is linearly related to the network average flow and can be approximated with ease, as illustrated in (Geroliminis and Daganzo, 2008; Zhou and Gayah, 2021). Then the intraregional completion rate can be further computed as $M_{ii}(t)=M_{i}(t)-M_{ij}(t),(i\neq j)$. For fair comparison, we offer two sets of states for testing under either idealized full observability or real-world limited observability.
Idealized full observability:
$s_{t}=[n_{ij}(t),\Delta n_{ij}(t),\Delta^{2}n_{ij}(t),M_{ij}(t)]$ (8)
Real-world limited observability:
$s_{t}=[n_{i}(t),\Delta n_{i}(t),\Delta^{2}n_{i}(t),M_{ij}(t)]$ (9)
@@
The action $a_{t}\in\mathcal{A}$ is defined the same as the control variables $u_{ij}(t)$ in Section 3. For the reward term $r_{t}$, while Zhou and Gayah (2021) uses merely the completion rate, in our proposed algorithm, the reward is defined with additional two terms, i.e., the damping term $r_{dam}(t)$ and the redundancy term $r_{red}(t)$, as Eq. 4 shows. When $[\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ or their $[\Delta n_{i}(t),\Delta^{2}n_{i}(t)]$ are incorporated into the state space $\mathcal{S}$ of the algorithm, significant oscillations in the perimeter control variables can often be observed, especially under scenarios with disruptions. As perimeter control is composed of coordinated traffic lights at the border of different regions, oscillating actions will result in fast varying green splits between consecutive cycles. Even though such oscillations may have a minor impact on traffic performance, the operation of signal lights in the real world should be as stable as possible. Therefore, a damping term $r_{dam}(t)$ is introduced into the reward function to penalize potential oscillatory actions. While the absolute difference between two consecutive actions is always below $1$, the constant $\xi_{1}$ bounds the maximal penalty while the exponent $\xi_{2}$ determines how fast the penalty decays when the change of control variables becomes small, indicating a large $\xi_{2}$ only penalizes the agent when the oscillation is substantial.
$r_{dam}(t)=-\xi_{1}\sum_{i,j\in\{1,2\}}|u_{i,j}(t)-u_{i,j}(t-1)|^{\xi_{2}},\quad i\neq j$ (10)
@@
The last term $r_{red}(t)$ in the objective function $J_{anti}$ in Eq. 4 acts as an additional term to build up redundancy in the system, resembling the buffer times in train scheduling (Corman et al., 2018). However, instead of a fixed buffer time commonly seen among most European railway networks (Jovanović et al., 2017), we create a dynamic redundancy term emphasizing the derivatives in a similar way as the differences of the regional accumulation, considering real-world limited observability. Here, we summarize $r_{red}(t)$ as the sum of two terms $r_{red}(t)=H(t)+\Delta H(t)$, with $H(t)$ being an overall term representing the first derivative and $\Delta H(t)$ representing the second derivative, which can be further expanded as:
@@
The binary variable $\alpha_{i}(t)$ in $H(t)$ was designed to reward the agent when moving towards the desired direction on the MFD. For instance, when the traffic state lies in the congested zone of the MFD, the gradient of any step will be negative. However, a penalty should be applied when vehicle accumulation gets larger, while rewarding the agent when accumulation decreases.
@@
Here, $f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})$ computes a reduction factor to constrain the impact of the $r_{red}(t)$ term when the accumulation is around the critical accumulation $n_{i,\rm crit}$, where the $r_{red}(t)$ term should have the greatest impact. A modified trigonometric function is applied to realize this purpose. Other functions, such as normal distribution, should also be valid for achieving the same goal.
@@
We illustrate the impact of $H(t)$ and $\Delta H(t)$ on building redundancy on the MFD in Fig. 4 and 5. The first derivative $H(t)$ in Fig. 4(a) rewards the agent when moving toward the critical accumulation $n_{i,\rm crit}$ to maximize trip completion. However, when $n$ approaches $n_{i,\rm crit}$, this term drops significantly and becomes a penalty when $n$ exceeds $n_{i,\rm crit}$. In Fig. 4, we showcase the influence of this term on the MFD. With increasing weight coefficient $\omega_{h}$, $n_{i,\rm crit}$ of the modified MFD becomes lower compared to the original, and the reward decreases faster after the accumulation exceeds $n_{i,\rm crit}$. In this way, redundant overcompensation has been established to prevent accumulation from exceeding the critical accumulation when disruption happens unexpectedly.
@@ 5 Experiment application
Following the contribution map illustrated in Fig. 1, we test our proposed antifragile perimeter control algorithm against three state-of-the-art perimeter control algorithms as benchmarks:
    MPC perimeter control proposed in Geroliminis et al. (2013).
    MPC-MHE perimeter control proposed in Sirmatel and Geroliminis (2020).
    Baseline RL perimeter control algorithm proposed in Zhou and Gayah (2021).
It is worth mentioning that this work employs MPC-MHE differently in comparison to Sirmatel and Geroliminis (2020), where sensor measurement noise is assumed and addressed by MHE through state estimation. In contrast, we presume the presence of disruption but perfect sensor measurements as the real states. As a result, since there is no need to estimate the system states, MPC-MHE does not provide additional benefits over MPC alone in the context of demand disruptions. However, for supply disruptions, we introduce a disruption magnitude reduction coefficient as an additional state variable, and MPC-MHE can be applied to estimate this coefficient and thereby the shift of the model.
Since transportation systems inherently manage the imbalance between demand and supply, real-world disruptions can also be broadly categorized as either a demand or a supply disruption. Therefore, the performance of the studied perimeter control algorithms in this paper is tested under demand or supply disruptions, as illustrated in Fig. 6. A demand disruption can be easily understood as a surging traffic demand due to social events or similar occurrences. Population growth and urbanization can also be considered a long-term and gradually incremental demand disruption. On the other side, a supply disruption can represent a link-level or network-level capacity drop due to adverse weather or major accidents, with network-level capacity drop being able to be reflected by the decrease of the MFD profile, as in Lu et al. (2024). Increasing traffic demand often leads to more frequent and severe traffic accidents (Dickerson et al., 2000). Other than the deterministically linearly increasing magnitude of disruptions, as represented by the solid linear red line, we also tested the algorithms of incremental magnitudes with uncertainties taken into account, as illustrated by the zigzag dotted red line, which will be explained in detail in the following section.
@@ 5.1 Simulation environment parametrization
We simulate a cordon-shaped urban network with inner and outer regions represented by different MFDs, as Fig. 2(b) shows. These MFDs are largely the same as in Zhou and Gayah (2021) and were originally approximated through the Yokohama loop detector dataset (Geroliminis and Daganzo, 2008). However, since the MFDs in Zhou and Gayah (2021) are formulated as piecewise functions and are, although continuous, not differentiable, the lack of differentiability within MFDs can cause fluctuations when computing the first and second differences and harm the efficacy of the redundancy term. Therefore, a minor modification has been made to slightly increase the gridlock accumulation by merely $3\%$, so that the whole MFD can be both continuous and differentiable. Other important indicators, e.g., critical accumulation and maximal trip completion, remain the same as in Zhou and Gayah (2021).
Traffic demand under no disruption is approximated based on Geroliminis and Daganzo (2008). As the real-world peak hour traffic demand profile has more resemblance to a Gaussian distribution instead of being a simple trapezoidal shape (Mazloumi et al., 2010), the base demand $q_{ij}(t)=[q_{11}(t),q_{12}(t),q_{21}(t),q_{22}(t)]$ is illustrated as the blue curves in Fig. 7(a). It consists of two components, which are a constant value $q_{ij,c}=[q_{11,c},q_{12,c},q_{21,c},q_{22,c}]=[0.2,0.4,0.1,0.3]$ in $\mathrm{veh/s}$ and a Gaussian term $q_{ij,N}(t)=C_{ij,N}\mathcal{N}_{ij}(\mu_{ij,N},\sigma_{ij,N})$ with total number of vehicles following Gaussian as $C_{ij,N}=[3000,10000,2000,7000]$ and the mean and variance of the distribution being $\mu_{ij,N}=[1800,1800,1800,1800]$ and $\sigma_{ij,N}=[1200,1500,900,1200]$ in seconds. Even though real-world disruptions can happen to both the outer and the inner regions, exerting a negative influence on the overall performance, we focus on the critical scenarios where the potential of perimeter control can be better reflected. As the inner region has a significantly smaller MFD profile compared to the outer region, hence, a surging disruptive traffic demand is assumed to be generated within the inner region. Therefore, when testing the performance of different algorithms under a growing demand disruption, $q_{22,N}$ is increased to simulate traffic from within the city center, reaching a total disruption demand of $12000$ vehicles at episode $75$, accounting for approximately one-third of the normal demand.
(a) Base demand profile and demand disruption
(b) MFD and supply disruption for the inner region
@@
Supply disruptions can manifest in various forms, such as adverse weather conditions, traffic accidents, and road maintenance. However, limited research has been dedicated to understanding the exact correlation between such events and their impact on the MFD. In this work, we adopt a proportional decrease in both jam density and maximal completion of the MFD, as shown in Fig. 7(b), in resemblance to Kim and Yeo (2017); Lu et al. (2024); Sun et al. (2024), representing lane closure events due to road construction works, traffic accidents, and natural disasters. Similar to testing the critical demand disruption on $q_{22,N}$, the critical scenario for supply disruption is the inner MFD being compromised while the outer MFD remains intact. Note that supply disruptions can be modeled in different ways for other events. For instance, considering the performance loss due to vehicle infrastructure interaction, Ambühl et al. (2020) proposed the concept of infrastructure potential $\lambda$ to represent how efficiently the network infrastructure is utilized, and a smaller $\lambda$ indicates the infrastructure is used with higher efficiency.
@@
The total simulation duration $T$ is 3 hours, with each time step $\Delta t$ taking $180$ seconds, and the third and last hour has little demand as it mainly serves as the unloading process. Another constraint is the lower and upper bounds for the perimeter control variable $u_{ij}(t)\in[0.1,0.9]$ to represent the real-world constraints from traffic signals (Geroliminis et al., 2013). The initial vehicle accumulation is set to be $n_{ij}(0)=[600,1300,300,2400]$ so that the accumulation remains approximately at an equilibrium at the beginning of the simulation. Each scenario is run $25$ times since randomness is inherent in RL, and its performance may vary across simulation iterations. Each iteration lasts for $75$ episodes, with the first $50$ under no additional disruption so that the RL agent can be properly trained under the base demand profile with predefined training noise. Incremental magnitude of disruptions is introduced in the subsequent $25$ episodes, with the value deliberately set to be the same as the number of simulation iterations. Under disruption uncertainties, as illustrated by the zigzag lines for the simulation scenarios in Fig. 6. A list of multipliers $\epsilon=[\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{24},\epsilon_{25}]$ following a normal distribution $\mathcal{N}\sim(1,0.15)$ is randomly generated to introduce uncertainties by computing the dot product with the list of disruption magnitudes. The same multiplier list is employed for all scenarios that are subject to disruption magnitude uncertainties. Also, it is shuffled by 1 for each simulation iteration. For example, the list would be $\epsilon=[\epsilon_{2},\epsilon_{3},\ldots,\epsilon_{25},\epsilon_{1}]$ for the second iteration. Through reshuffling, the disruption magnitudes from different episodes will experience exactly the same list of uncertainty multipliers but with different sequences.
@@
The most important hyperparameters for both the baseline and the proposed RL algorithms should be the same and are summarized in Table 2. Note that the minimal learning rates and noise scale are not set to be a rather small value as common RL algorithms do, we aim to get a trade-off between optimality and adaptiveness when the algorithm is experiencing disruptions, which can be demonstrated by the results for both superior performance and antifragility in Section 6. The coefficients $\xi_{1}$ and $\xi_{2}$ for the damping term $r_{dam}(t)$ in Eq. 10 to counter the potential oscillations actions are set to be $1$ and $6$. The weights $\omega_{h}$ and $\omega_{\Delta h}$ for the redundancy term $r_{red}(t)$ is set to be $0.1$ and $0.2$. The number of simulations generated under the Ape-X architecture is 32 per training episode. Both the prediction horizon and control horizon of MPC and MPC-MHE are set to 10.
@@
Learning rate decay 0.98 0.98